OpenStudy (anonymous):

Prove that the fifth of root of 15 is irrational

OpenStudy (anonymous):

same proof by contradiction as \(\sqrt2\) is irrational suppose it is, and derive a contradiction \[\sqrt[5]{15}=\frac{a}{b}\\15=\frac{a^5}{b^5}\\15b^5=a^5\] and contradiction like with prime factorization on both sides of the equal sign

OpenStudy (anonymous):

Can you go into more detail about last part?

OpenStudy (anonymous):

I follow up until the last part. I've done the proof of 2^1/2, but I dont understand how to compare it to any other root

OpenStudy (anonymous):

yes i can be in more detail but i will say it in english you can write the math

OpenStudy (anonymous):

on the right hand side you have \(a^5\) however \(a\) factors as the product of primes, when you raise it to the 5th power, each prime will have a power that is a multiple of 5

OpenStudy (anonymous):

on the left hand side you have \(b^5\) and the same argument holds, each prime factor will be raised to a multiple of 5 for that but the left hand side is \(15b^5=3\times 5\times b^5\) so the prime \(3\) will have an exponent that is NOT a multiple of 5, as will the prime \(5\) since there is only one way to factor a number as the product of primes (fundamental theorem of arithmetic) that is a contradiction

OpenStudy (anonymous):

okay, i don't really understand that, but it works. Thank you